OpenStudy (anonymous):

Find the derivative of f(x) = 7 / x at x = 1.

jimthompson5910 (jim_thompson5910):

f(x) = 7/x f(x) = 7x^(-1) f ' (x) = 7(-1)*x^(-2) f ' (x) = -7*x^(-2) now plug in x = 1 and evaluate